OpenStudy (anonymous):

Simplify each radical expression? √7 ( √14 +√3)

OpenStudy (ivancsc1996):

\[\sqrt{7}\sqrt{14}+\sqrt{7}\sqrt{3}=\sqrt{7 \times 14}+\sqrt{7\times3}=\sqrt{98}+\sqrt{21}=7\sqrt{2}+\sqrt{21}\]

OpenStudy (anonymous):

Thank you so much, just what i was looking for! @ivancsc1996

OpenStudy (ivancsc1996):

No problem :)